# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v2

@@ 7.5. Computational cost analysis
QPP Method Inference latency per query (ms)
1 2 3 4
@@
Table 7 shows the online QPP latency of QPP-GenRE integrated with LLaMA-7B and other BERT-based supervised QPP baselines, on TREC-DL 19, on a single NVIDIA A100 GPU. We compute the inference latency when queries are processed individually. For QPP-GenRE, we consider judging depths at 10, 100, and 200; QPP-GenRE can use batch acceleration for judging items for the same query because each item in a ranked list for a query is independent of each other.111111qppBERT-PL first splits a ranked list with 100 items into 25 chunks and then predicts the number of relevant items in each chunk. For a fair comparison, we put 25 chunks into one batch for acceleration. Although QPP-GenRE is more expensive than all baselines when predicting one measure due to the much larger parameter size of LLaMA-7B compared to BERT, QPP-GenRE has lower latency compared to some baselines when predicting multiple IR evaluation measures because multiple measures can be derived from the same set of relevance judgments at no additional cost. E.g., while QPP-GenRE is 56% more expensive than M-QPPF for predicting one measure, it becomes more efficient when predicting 2 or more metrics than M-QPPF. Nevertheless, we acknowledge that QPP-GenRE has higher computational costs than supervised QPP methods when predicting a single measure. Conversely, regression-based QPP baselines (NQA-QPP, BERTQPP and M-QPPF) need to train separate models for different IR evaluation metrics. Although qppBERT-PL is not optimized to learn to output one specific IR evaluation measure, qppBERT-PL does not achieve a promising QPP quality (see Table 2 and 4).
@@
We argue that QPP-GenRE’s latency is still much smaller than some high-performing LLM-based re-rankers. E.g., Sun et al. (2023b) show that a GPT-4-based listwise re-ranker needs 10 API calls (one call takes 3,200ms) to re-rank 100 items for a query, resulting in 32,000ms in total, which is around 20 times worse than QPP-GenRE’s latency with a judging depth of 100. QPP-GenRE can well fit some knowledge-intensive professional search scenarios where QPP quality is prioritized or users may have a higher tolerance level for latency, such as patent search (Lupu et al., 2013), legal search (Tomlinson et al., 2007). Besides using QPP online, QPP can also be used to analyze a search system’s performance in an offline setting (Faggioli et al., 2023d).
@@ 8. Conclusions & Future Work
We have proposed a new QPP framework, QPP-GenRE, which models QPP from the perspective of predicting IR evaluation measures based on automatically generated relevance judgments. We have devised an approximation strategy for predicting an IR evaluation measure considering recall, which only judges a limited number of items in a given ranked list for a query, to avoid the cost of traversing the entire corpus to find all relevant items; the approximation strategy also enables us to study into the impact of various judging depths on QPP quality. We have explored using open-source LLMs for generating relevance judgments, to ensure scientific reproducibility. In addition, we have examinged training open-source LLMs with parameter-efficient fine-tuning (PEFT) on human-labeled relevance judgments, to improve the quality of relevance judgment generation and QPP.
@@
Experiments on datasets from the TREC-DL 19–22 tracks demonstrate that QPP-GenRE significantly surpasses existing QPP approaches, achieving state-of-the-art QPP quality in assessing lexical and neural rankers for either a precision-oriented IR metric or an IR metric considering recall. Moreover, we have shown that QPP-GenRE has the potential to conduct QPP more accurately when integrated with a more effective LLM, has a good compatibility with other types of relevance prediction methods (e.g., an LLMs-based re-ranker), and exhibits good interpretability.
@@
QPP-GenRE has the potential to facilitate the practical use of QPP. The limited accuracy and interpretability of current QPP methods make them difficult to use in practical applications (Arabzadeh et al., 2024). However, QPP-GenRE demonstrates significantly improved QPP accuracy and better interpretability, enhancing the reliability of QPP results and potentially facilitating the practical use of QPP. Especially, QPP-GenRE has the potential to benefit some knowledge-intensive professional search scenarios, e.g., legal (Tomlinson et al., 2007) or patent search (Lupu et al., 2013). In such scenarios, accurate QPP is prioritized, interpretable QPP results are needed, and users may have a higher tolerance level for latency. QPP-GenRE also has the potential for practical application in commercial search engines: commercial search engines receive many frequent and repeated queries, and QPP-GenRE can improve QPP efficiency by reusing stored relevance judgments for repeated query-item pairs and only generating relevance judgments for new query-item pairs. Moreover, QPP-GenRE can be used to analyze the ranking quality of a search system in an purely offline setting (Faggioli et al., 2023d), where latency is not necessarily an issue.
@@
Second, QPP-GenRE is a QPP framework that can be integrated with various relevance prediction approaches. We show the success of QPP-GenRE equipped with LLaMA-7B, Llama-3-8B, Llama-3-8B-Instruct as well as a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a). Exploring various LLMs to find the optimal one for relevance prediction is beyond the scope of our work. However, in future, we believe it is valuable to investigate QPP-GenRE’s performance integrated with other open-source LLMs (e.g., Mistral (Jiang et al., 2023)) as relevance judgment generators. It is also interesting to adapt pairwise or listwise LLM-based re-rankers into relevance judgment generators and integrate QPP-GenRE with them.
@@
Fourth, while QPP-GenRE exhibits a promising QPP quality and can be used in scenarios where QPP quality is prioritized and users have a higher tolerance level for latency, e.g., patent search or post analysis, it is worth improving QPP-GenRE’s efficiency in future to widen its scope of applications. We plan to investigate (i) the use of multiple GPUs because judging each item in a ranked list is independent of each other, (ii) distilling knowledge from LLMs to smaller language models (Gu et al., 2023), and (iii) compressing LLMs by using lower-bit (e.g., 2-bit) quantization (Chee et al., 2023) or using low-rank factorization (Xu et al., 2023).
@@ Acknowledgements.
This research was partially supported by the China Scholarship Council (CSC) under grant number 202106220041, the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), DPG Media, RTL, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v3

@@ 7.6. Computational cost analysis
Inference latency per query (ms)
QPP Method 1 2 3 4
@@
Table 11 shows the online QPP latency of QPP-GenRE integrated with LLaMA-7B and other BERT-based supervised QPP baselines, on TREC-DL 19, on a single NVIDIA A100 GPU. We compute the inference latency when queries are processed individually. For QPP-GenRE, we consider judging depths at 10, 100, and 200; QPP-GenRE can use batch acceleration for judging items for the same query because each item in a ranked list for a query is independent of each other.111111qppBERT-PL first splits a ranked list with 100 items into 25 chunks and then predicts the number of relevant items in each chunk. For a fair comparison, we put 25 chunks into one batch for acceleration. Although QPP-GenRE is more expensive than all baselines when predicting one measure due to the much larger parameter size of LLaMA-7B compared to BERT, QPP-GenRE has lower latency compared to some baselines when predicting multiple IR evaluation measures because multiple measures can be derived from the same set of relevance judgments at no additional cost. E.g., while QPP-GenRE is 56% more expensive than M-QPPF for predicting one measure, it becomes more efficient when predicting 2 or more metrics than M-QPPF. Nevertheless, we acknowledge that QPP-GenRE has higher computational costs than supervised QPP methods when predicting a single measure. Conversely, regression-based QPP baselines (NQA-QPP, BERTQPP and M-QPPF) need to train separate models for different IR evaluation metrics. Although qppBERT-PL is not optimized to learn to output one specific IR evaluation measure, qppBERT-PL does not achieve a promising QPP quality (see Sections 6.1 and 6.2).
@@
We argue that QPP-GenRE’s latency is still much smaller than some high-performing LLM-based re-rankers. E.g., Sun et al. (2023b) show that a GPT-4-based listwise re-ranker needs 10 API calls (one call takes 3,200ms) to re-rank 100 items for a query, resulting in 32,000ms in total, which is around 20 times worse than QPP-GenRE’s latency with a judging depth of 100. QPP-GenRE can well fit some knowledge-intensive professional search scenarios where QPP quality is prioritized or users may have a higher tolerance level for latency. Besides using QPP online, QPP can also be used to analyze a search system’s performance in an offline setting (Faggioli et al., 2023d).
(a) TREC-DL 19
(b) TREC-DL 20
Figure 8. The average number of actual relevance predictions required for sequentially predicting the performance of BM25, ANCE, and TAS-B using the relevance judgment caching mechanism, with a judging depth of $n$ equal to 10 or 100, on TREC-DL 19 and 20.
Lastly, in order to enhance the efficiency of QPP-GenRE, we propose a relevance judgment caching mechanism. It reuses previously predicted relevance judgments for the same query when predicting the performance of new rankers. As a result, this mechanism helps conserve computational resources by eliminating the need to recompute relevance judgments that are shared among multiple rankers. Figure 8 shows the number of actual relevance predictions required for sequentially predicting the performance of BM25, ANCE, and TAS-B using the relevance judgment caching mechanism, with a judging depth of $n$ equal to 10 or 100, on TREC-DL 19 and 20. We found that our proposed relevance judgment caching mechanism can reduce the number of LLM calls for relevance prediction by approximately 30%. For instance, on TREC-DL 19, with a judging depth of 10, the caching mechanism results in 21.15 LLM calls on average when sequentially predicting the performance of the three rankers (10 for BM25, 7.06 for ANCE, and 4.09 for TAS-B). In contrast, without using this mechanism, 30 LLM calls would be required (3 $\times$ 10).
@@ 8. Conclusions & Future Work
We have proposed a new QPP framework, QPP-GenRE, which models QPP from the perspective of predicting IR evaluation measures based on automatically generated relevance judgments. We have devised an approximation strategy for predicting an IR evaluation measure considering recall, which only judges a limited number of items in a given ranked list for a query, to avoid the cost of traversing the entire corpus to find all relevant items; the approximation strategy also enables us to study into the impact of various judging depths on QPP quality. We have explored using open-source LLMs for generating relevance judgments, to ensure scientific reproducibility. In addition, we have examined training open-source LLMs with parameter-efficient fine-tuning (PEFT) on human-labeled relevance judgments, to improve the quality of relevance judgment generation and QPP.
@@
Experiments on datasets from the TREC-DL 19–22 tracks demonstrate that QPP-GenRE significantly surpasses existing QPP approaches, achieving state-of-the-art QPP quality in assessing lexical and neural rankers for either a precision-oriented IR metric or an IR metric considering recall. Moreover, we have shown that fine-tuning open-source LLMs on human-labeled relevance judgments is crucial for obtaining reliable relevance prediction and QPP results. Fine-tuning much smaller LLMs for relevance judgment prediction can yield more effective relevance prediction and QPP than few-shot prompting with much larger models. In particular, a fine-tuned 3B model (Llama-3.2-3B-Instruct) offers the best trade-off between QPP quality and computational efficiency: it significantly outperforms 70B few-shot models and delivers performance comparable to fine-tuned 7B and 8B models. It implies that fine-tuning can offer strong performance at low inference costs. Furthermore, QPP-GenRE has the potential to conduct QPP more accurately when integrated with a more effective LLM, has a good compatibility with other types of relevance prediction methods (e.g., an LLMs-based re-ranker). Additionally, We have demonstrated that QPP-GenRE has great generalizability to the conversational search scenario. We have shown that QPP-GenRE exhibits good interpretability. Finally, we have found that our proposed relevance judgment caching mechanism can reduce LLM calls for relevance prediction by about 30%.
@@
QPP-GenRE has the potential to facilitate the practical use of QPP. The limited accuracy and interpretability of current QPP methods make them difficult to use in practical applications (Arabzadeh et al., 2024a). However, QPP-GenRE demonstrates significantly improved QPP accuracy and better interpretability, enhancing the reliability of QPP results and potentially facilitating the practical use of QPP. Especially, QPP-GenRE has the potential to benefit some knowledge-intensive professional search scenarios. In such scenarios, accurate QPP is prioritized, interpretable QPP results are needed, and users may have a higher tolerance level for latency. QPP-GenRE also has the potential for practical application in commercial search engines: commercial search engines receive many frequent and repeated queries, and QPP-GenRE can improve QPP efficiency by reusing stored relevance judgments for repeated query-item pairs and only generating relevance judgments for new query-item pairs. Moreover, QPP-GenRE can be used to analyze the ranking quality of a search system in an purely offline setting (Faggioli et al., 2023d), where latency is not necessarily an issue.
@@
Second, QPP-GenRE is a QPP framework that can be integrated with various relevance prediction approaches. We show the success of QPP-GenRE equipped with various open-source LLMs as well as a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a). Exploring various LLMs to find the optimal one for relevance prediction is beyond the scope of our work. However, in future, we believe it is valuable to investigate QPP-GenRE’s performance integrated with other open-source LLMs as relevance judgment generators. It is also interesting to adapt pairwise or listwise LLM-based re-rankers into relevance judgment generators and integrate QPP-GenRE with them.
@@
Fourth, while QPP-GenRE exhibits a promising QPP quality and can be used in scenarios where QPP quality is prioritized and users have a higher tolerance level for latency, e.g., patent search or post analysis, it is worth improving QPP-GenRE’s efficiency in future to widen its scope of applications. We plan to investigate (i) the use of multiple GPUs because judging each item in a ranked list is independent of each other, (ii) distilling knowledge from LLMs to smaller language models (Gu et al., 2023), (iii) compressing LLMs by using lower-bit (e.g., 2-bit) quantization (Chee et al., 2023) or using low-rank factorization (Xu et al., 2023), and (iv) proposing an adaptive sampling approach that selects only a subset of documents from a ranked list for LLM-based relevance prediction to optimize the trade-off between judgment cost and QPP performance.
@@ Acknowledgements.
We thank our reviewers and associate editor for their constructive feedback and suggestions. This research was partially supported by the China Scholarship Council (CSC) under grant number 202106220041, the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21, which is (partly) financed by the Dutch Research Council (NWO), project ROBUST with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO) and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023, and the FINDHR (Fairness and Intersectional Non-Discrimination in Human Recommendation) project that received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.